OpenStudy (anonymous):

can someone please help me ive been stuck on this for a month and im so behind :(

OpenStudy (isaiah.feynman):

Are you ready?

OpenStudy (anonymous):

yes

OpenStudy (isaiah.feynman):

For number 4, we would need to find AC first, because we can use it to find the angles we need.

OpenStudy (isaiah.feynman):

Are you familiar with the Pythagorean theorem?

OpenStudy (anonymous):

not really... all ive been seeing is a^2+b^2=C^2

OpenStudy (isaiah.feynman):

Here,  For any right triangle, the sum of the square of the opposite and adjacent sides equals the square of the hypotenuse.

OpenStudy (anonymous):

ok so for my problem how do i plug it in ? 10+4=c?

OpenStudy (isaiah.feynman):

In your problem, we have

OpenStudy (isaiah.feynman):



OpenStudy (isaiah.feynman):

Can you see that AC corresponds to "a" in the example I showed?

OpenStudy (anonymous):

what do you mean when you put corresponds

OpenStudy (isaiah.feynman):

In my example "a" is the lower side of the triangle, in your problem, AC is also the lower part of the triangle. Same sides.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i see that

OpenStudy (isaiah.feynman):

Good. Since its that way we can solve for AC. Since \[c^{2} = a^{2} + b^{2} \] We can solve for a to give us \[a = \sqrt{c^{2} - b^{2}}\] That also means we can solve for AC to give us \[AC^{2} = \sqrt{10^{2} - 4^{2}}\]

OpenStudy (anonymous):

oh ok so then do i solve the last equation?

OpenStudy (isaiah.feynman):

Yes. Sorry I made a mistake, it should be AC NOT AC^2

OpenStudy (anonymous):

ok so how do i do that minus 10^2-4^2 and that gives me 84

OpenStudy (anonymous):

\[\sqrt{84}\]

OpenStudy (anonymous):

and do i solve that?

OpenStudy (isaiah.feynman):

Yes.

OpenStudy (anonymous):

\[2\sqrt{21}\]

OpenStudy (anonymous):

or 9.1651513

OpenStudy (isaiah.feynman):

Smart!

OpenStudy (anonymous):

thank you :) so do i do that for all of them and can you help me with another one

OpenStudy (isaiah.feynman):

Now we have figured out AC. Time to get angles!

OpenStudy (anonymous):

oh yeah lol almost forgot i wasnt done

OpenStudy (isaiah.feynman):

So do you know trig ratios?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

so 10= 4/c

OpenStudy (isaiah.feynman):

Oh I forgot something,

OpenStudy (anonymous):

oh idk how to do that

OpenStudy (isaiah.feynman):

Okay, a little teaching on trig ratios. In a right triangle we have the following properties,

OpenStudy (anonymous):

so sin(a)= 4/10

OpenStudy (isaiah.feynman):

Excellent!

OpenStudy (anonymous):

what do we do after that

OpenStudy (isaiah.feynman):

To find A itself we need to introduce something else, its called the inverse sine. It works this way \[\sin \theta =\frac{ a }{ c }\] \[\theta = \sin^{-1} \frac{ a }{ c }\]

OpenStudy (isaiah.feynman):

I did this \[\sin^{-1} (\sin \theta) = \sin^{-1} \frac{ a }{ c }\] Which then gave me \[\theta = \sin^{-1} \frac{ a }{ c }\]

OpenStudy (anonymous):

do i plus a in or do we just leave it like that

OpenStudy (isaiah.feynman):

Because of your question you should plug in a and c.

OpenStudy (anonymous):

ok so it would be \[\Theta= \sin -1 4/10 \]

OpenStudy (anonymous):

opps o forgot to space

OpenStudy (anonymous):

but is that negative 1 a exponent or a actual -1

OpenStudy (isaiah.feynman):

Its just a symbol. It means nothing.

OpenStudy (anonymous):

oh ok

OpenStudy (isaiah.feynman):

So it should be \[\sin^{-1} \frac{ 4 }{ 10 } = \sin^{-1} \frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

how do i solve that

OpenStudy (isaiah.feynman):

Use your calculator. Press shift + sin first, then 2/5, then =

OpenStudy (anonymous):

i got .4

OpenStudy (isaiah.feynman):

Nope. Try again. It should be 23.5 degrees.

OpenStudy (anonymous):

ohhh i got it

OpenStudy (anonymous):

and thats for A

OpenStudy (isaiah.feynman):

To find B. We do the same thing. The side opposite B is 9.1 (Remember this is AC we solved for), and the hypotenuse is c = 10. So solve for B using what I showed.

OpenStudy (anonymous):

is b 66.42 degrees

OpenStudy (isaiah.feynman):

Yes.

OpenStudy (anonymous):

yayyyyy :)

OpenStudy (isaiah.feynman):

You learn quick.

OpenStudy (anonymous):

only because you actually helped me :)

OpenStudy (isaiah.feynman):

:)

OpenStudy (isaiah.feynman):

Can you solve number 5 using all you've learned so far?

OpenStudy (anonymous):

yes i think i can finish the rest by myself thanks for the help! you made my day :)))))

OpenStudy (isaiah.feynman):

Good stuff!